OpenStudy (anonymous):

The does anyone know the formula for finding someones target heart rate for exercise

OpenStudy (anonymous):

Target heart rate zones range anywhere from 50% to 100% of your maximum heart rate (your maximum heart rate is based on your age). Aerobic exercise is anything less than 85%, and anaerobic exercise is anything above that. A nice starting point for a sedentary individual is somewhere in the range from 50% to 65% (you can always increase as you get more fit) and 65% to 85% for more conditioned individuals. I recommend the heart-rate reserve method for calculating a target heart rate. Here's the formula and an example of the method for someone 27 years of age, assuming a resting heart rate of 70 bpm, and a training range of 70%. If you plug in other values, you can get other ranges. 220-age = Max HR. Subtract resting heart rate from Max HR = Heart Rate Reserve (HRR). Multiply HRR times percent you want to train at. Add back resting heart rate. Assuming a resting heart rate of 70 bpm, 27 years old, and 70% training range: 220 - 27 = 193 193 - 70 = 123 123 x .70% = 86 86 + 70 = 156 Please note: There's been some recent research to suggest a new way of estimating maximum heart rate. The formula is the following: Multiply 0.7 times your age . Subtract that number from 208. An example if you're 26 years old is: 0.7 x 26 = 18, then 208 - 18 = 190. You'd then take 190 and plug it in as usual to the formula above. This new formula makes a slightly bigger difference as you get older.

OpenStudy (anonymous):

@shann_kwee